Decide how the electrons in the following table would be listed, if they were listed in order of increasing electrostatic potential energy. That is, select 1 next to the electron with the lowest electrostatic potential energy, select 2 next to the electron with the second lowest electrostatic potential energy, and so on. electrostatic potential energy electrorn location A 3.4 nm from a cation with a charge of +2(choose one) B 3.4 nm from a cation with a charge of +3(choose one) C2.1 nm from a cation with a charge of +3 (choose one) D 4.8 nm from a cation with a charge of +2 (choose one)
Answer #1

Electrostatic potential energy is given as:

E = kQar where k = 9*109 N.m2/C2

Charge on electron is -1. So Q = -1.

From above equation we can see that for same radius potential energy is directly proportional to charge. So B<A (Remember Q = -1, so sign is negative)

For same charge, lesser the radius more is potential energy. so C<B and similarly A<D

From above relations we can say that C<B<A<D

For exact values we can calculate as follows,

A: k*-1*2/(3.4)2 = -0.173k

B: k*-1*3/(3.4)2 = -0.259k

C: k*-1*3/(2.1)2 = -0.680k

D: k*-1*2/(4.8)2 = -0.086k

From these values also we get the order of potential energy as D>A>B>C